Given The Cobb- Douglas Production Function For Preshzo Company Limited As Follows: đť‘„=K0.25L0.75 She Estimates That The Cost Of Labour Is $3 And The Cost Of Capital $4. You Are Required To Do The Following: A) If She Wants To Produce 50kg Of Are Output, Formulate Both The Objective And Constraint Functions For Preshzo Company Limited. B) Find The Minimum

1
Expert's answer
2021-10-10T17:59:36-0400

Solution:

A.). The objective and constraint functions are as follows:

Constraint function: K0.25L0.75 = 50

Objective function: C = wL + rK


B.). To derive the minimum cost of labor and capital required to produce 50kg units of output:

MRTS = "\frac{MP_{L} }{MP_{K}} = \frac{w}{r}"


MPL = "\frac{\partial Q} {\partial L} =" 0.750.25L-0.25


MPK = "\frac{\partial Q} {\partial K} =" 0.25K-0.75L0.75


MRTS = "\frac{MP_{L} }{MP_{K}} = \frac{0.75^{0.25} L^{-0.25} }{0.25K^{-0.75} L^{0.75} } = \frac{3K}{L}"


"\frac{3K}{L} = \frac{w}{r} = \frac{3}{4}"


"\frac{3K}{L} = \frac{3}{4}"

L = 4K


Substitute in the Cobb-Douglas function to derive Capital:


50 = K0.25L0.75 = K0.254K0.75

K = 12.5

L = 4K = 12.5 "\times" 4 = 50

The minimum cost of labor = 50

The minimum cost of capital = 12.5